.05x+.25y=11 and .15x+.05y=-12
Answer by stanbon(75887)

.05x+.25y=11
.15x+.05y=-12
--------------------
Multiply thru both equations by 100.
--------------------
5x + 25y = 1100
15x + 5y = -1200
-----------------------
Multiply thru the 1st equation by 3:
15x + 75y = 3300
15x + 5y = -1200
-------------------
Substract the 2nd from the 1st:
70y = 4500
y = 64.29
------
Substitute into 5x + 25y = 1100
5x + 25*64.29 = 1100
5x = -507.14
x = -101.42
==============
Check in 15x+5y = -1200
15(-101.42) + 5(64.29) = -1199.85
Answers check.
================
Cheers,
Stan H.